The values are in direct variation. Find the missing value.
(3, 2) and (6, y)

4
5
1
12



The values are in direct variation. Find the missing value.
(x, 8) and (6, -16)
-21.3
-3
3
-12

Answer by ewatrrr(24785)

y = kx
2 = 3k
2/3 = k
y = (2/3)x , x+=+6, y = 4
.........
y = kx
-16 = 6k
-16/6 = k = -8/3
y = (-8/3)x, y+=+8 , (-3/8)(8) = x = -3